William opened two investment accounts for his grandson’s college fund. The first year, these investments, which totaled $18,000, yielded $831 in simple interest. Part of the money was invested at 5.5% and the rest at 4%. How much was invested at each rate?
Answer by mananth(16946)

Investment Part I 5.50% per annum ----x
Investment part II 4.00% per annum ----y
The sum of the investments is $18,000.00
The sum of individual interests = $831.00
x+y 18000 ------------------------1
5.50%x +4.00%y = $831.00
Multiply by 100
5.5x + 4y = $83,100.00 --------2

Multiply (1) by -5.5
we get

-5.5x -5.5y =-99000.00

Add this to (2)

0x -1.5y =-$15,900.00

divide by-1.5

y = $10,600.00 investment at 4.00%
Balance $7,400.00 investment at 5.50%
CHECK
$7,400.00 @ 5.50% $407.00
$10,600.00 @ 4.00% $424.00
Total -------------------- $831.00